Find the domain and range
f(x)= 2- x^2

Found 2 solutions by edjones, stanbon:
Answer by edjones(8007)

the domain is all real numbers
the range is all real numbers =<2 or f(x)(-infinity, 2]
Ed

Answer by stanbon(75887)

Find the domain and range
f(x)= 2- x^2
Domain: all Real Numbers
Range : Since x^2 >=0 2-x^2 <=2
So, Range is "all Real Numbers <= 2"
===============================
Cheers,
Stan H.